OpenStudy (anonymous):

A group of 3 boys and 2 girls are seated in a row of 5 chairs find the probability that they will be seated alternately..plz help..

Parth (parthkohli):

You just have to find the number of ways they can sit, then find the number of ways they can sit alternately.

OpenStudy (artfis):

5! = 5 x 4 x 3 x 2 x 1 = 120 ways

Parth (parthkohli):

And the number of ways in which they can sit alternately is \(3 \times 2 \times 2 \times 1 \times 1\)

OpenStudy (anonymous):

3x2x2x1x1

OpenStudy (anonymous):

so??what is the answer?

Parth (parthkohli):

Number of ways they can sit alternately divided by the total number of ways.

OpenStudy (artfis):

@ParthKohli is right. My fault. Looked at the problem a lil bit different. My fault.

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

3!2! divided by 5!??is that correct??

OpenStudy (anonymous):

24

Parth (parthkohli):

yes, that is right.

OpenStudy (anonymous):

3C1x2C1x2C1=12...i think its 12...

Parth (parthkohli):

yes.

Parth (parthkohli):

12/120 = ...

OpenStudy (anonymous):

1/10

Parth (parthkohli):

Yeah!